# Racka: Efficient Hungarian LLM Adaptation on Academic Infrastructure

### v1

@@ 4.3 Packed Continual Pretraining with LoRA
To enable parameter-efficient continual pretraining, we employed Low-Rank Adaptation (LoRA) adapters. LoRA modules were inserted into all linear layers of the Qwen3-4B backbone. The original backbone weights remained frozen throughout training. The embedding and language modeling head layers are tied via transposed weight replication as in the Qwen3-4B base.
Our choice of LoRA hyperparameters was guided by recent best practices and empirical validation. Specifically, we set the LoRA rank ($r$) to $64$ and the scaling factor ($\alpha$) to $128$, following the heuristic of setting $\alpha$ to approximately twice the rank to ensure sufficient adaptation capacity. A dropout of $0.1$ was applied to the LoRA adapters to improve generalization. With this setup we train 0.520.52B parameters accounting for $12.5\%$ of the base model size. We used separate learning rates for LoRA and non-LoRA parameters: $1\times 10^{-4}$ for LoRA and $5\times 10^{-5}$ for non-LoRA parameters, both with a weight decay of $0.005$. Optimization was performed using AdamW. The learning rate schedule included a linear warmup phase covering $1\%$ of total training steps, followed by linear decay to zero which is shown to be beneficial (bergsma2025straightzerolinearlydecaying). We did not employ a specialized tokenizer healing phase.
To maximize training efficiency, we used input sequence packing (ding2024fewertruncationsimprovelanguage), merging shorter sequences into a single context window of up to 4096 tokens. Both the beginning-of-sequence and end-of-sequence delimiters were set to <|endoftext|>, as defined by the Qwen tokenizer. Packing was performed using a linear scaling algorithm prior to training. For computational efficiency, we used a standard full causal attention mask (as in (deepseekai2025deepseekv3technicalreport)), rather than a multi-segment lower-triangular mask, which would otherwise require quadratic memory and computational overhead (poedator4dmasks2024).
@@ 4.4 Training Infrastructure and Parallelization Strategy
We conducted performance measurements on a distributed setup consisting of two nodes, each equipped with four GPUs. In DDP configuration, the training throughput reached 2.92.93.03.0 seconds per iteration. In contrast, FSDP configuration delivered substantially lower performance, reaching 5.05.05.15.1 seconds per iteration. Mixed-precision training with the standard FSDP implementation is not possible, training with full precision resulted in OOM for our minimal training context length of 44k tokens. As DDP yielded superior scaling and GPU utilization, so we adopted it for all experiments as a hardware-aware choice for parallelization on the available infrastructure.
@@
4.5 Evaluation Protocol
Our evaluation protocol was designed to assess the model from two perspectives: its intrinsic ability to model language distributions and its practical capabilities on different Hungarian language model benchmarks. In this section we will introduce the intrinsic evalution and the mentioned benchmarks one-by-one.
Intrinsic Evaluation: We use perplexity (PPL) as our primary intrinsic metric. Perplexity measures the model’s uncertainty in predicting the next token in a sequence; a lower perplexity score indicates a better fit to the data distribution. (10.1121/1.2016299) We calculated perplexity on the held-out validation set containing a mix of Hungarian, English, German and code to monitor learning progress and assess for catastrophic forgetting.
LM-Eval-Harness benchmark: To measure the model’s practical utility without task-specific training, we conducted few-shot evaluations on a selection of Hungarian language understanding tasks (eval-harness). For this we utilized the lm-eval-harness with Hungarian translated benchmarks. We used the already implemented benchmarks ARC_HU, MMLU_HU, Hellaswag_HU and TruthfulQA_HU from lm-eval-harness, which is the work of (okapi). Additionally, we manually added the hungarian translated version of GSM8K (cobbe2021gsm8k), which is a mathematical reasoning dataset and was transletd by (thellmann2024crosslingual), to check the model reasoning capabilities on mathemical problems.
We ran all benchmarks separately, using the default config parameters defined in the original lm-eval-harness repository [^2] https://github.com/EleutherAI/lm-evaluation-harness. Table 3 shows our results.
HULU Benchmark The HULU benchmark offers a comprehensive set of supervised tasks covering a wide range of linguistic tasks, including acceptability, causal reasoning, sentiment classification, pronoun resolution, factual commitment, and reading comprehension. For the evaluation, we separately fine-tuned our model on each HULU subtask using a forked version of the benchmark repository [^3] https://github.com/nytud/HuLU. Fine-tuning was performed using full finetune and Low-Rank Adaptation (LoRA) matrices with the following key parameters: lora_rank=64lora\_rank=64, lora_alpha=128lora\_alpha=128, lora_dropout=0.1lora\_dropout=0.1, and batch_size=2batch\_size=2.
OpenHuEval benchmark To further evaluate the ability of the model to perform Hungarian reading comprehension and generation tasks, we conducted fine-tuned downstream evaluations using the OpenHuEval benchmark suite (openhueval). This collection focuses on cloze-style question answering and comprehension tasks designed to capture various aspects of contextual understanding, reasoning, and lexical inference in Hungarian. OpenHuEval comprises five datasets HuWildBench (real Hungarian user queries), HuSimpleQA (fact-based Hungarian questions), HuProverbRea (reasoning with Hungarian proverbs), HuMatchingFIB (option-based fill-in-the-blank), and HuStandardFIB (free-form fill-in-the-blank). The evaluation was made using the original repository of the benchmark [^4] https://github.com/opendatalab/OpenHuEval.
For comparison we evaluated the original Qwen3-4B, the Qwen3-4B-Base and the PULI-LlumiX-Llama-3.1 models on all benchmarks to check the development of the model after the training and to compare it with another hungarian LLM. The evaluation of the models were on a single A100 GPU for all benchmarks.
5 Results and Analysis
5.1 Tokenizer Performance
The primary goal of our tokenizer adaptation was to improve the encoding efficiency for Hungarian without degrading performance on the high-resource languages present in the base model. We evaluated this using subword fertility, which is the average number of tokens required to represent a single whitespace delimited word.
The results measured on the validation documents, presented in Table 2, confirm the success of our adaptation strategy. For Hungarian, the Racka tokenizer demonstrates a substantial improvement over the original Qwen-3 tokenizer. Subword fertility was reduced by over $46\%$, from $3.13$ down to $1.66$. This means that, on average, the same Hungarian text can be represented with almost half of the original Qwen tokens, leading to direct improvements in inference efficiency and latency. Crucially, this gain did not come at the same expense of performance on other languages with slight increase for German and a less than $24\%$ fertility increase for English. This indicates that our vocabulary extension successfully preserved the efficiency for high-resource languages.
Table 2: Tokenizer performance on the validation set: Subword fertility (lower is better) and relative change from Qwen-3 4B.
Language Qwen-3 4B Racka-4B Change (%)
@@ 5.1 Tokenizer Performance
5.2 Continual Pretraining Results
Following the successful tokenizer adaptation, we successfully conducted the final continual pretraining phase on the full 160B token dataset. The training ran for 326 357326\,357 steps on 64 GPUs in parallel, with a per-device batch size of $2$ and gradient accumulation set to $4$ resulting in an effective logical batch size of $512$ sequences with a context length of $4096$ for each model update. The training loss curve is depicted in 1 Figure. We see that even though the training loss begins to flatten after 50 00050\,000 steps, the validation perplexity continues to decrease, indicating that overfitting is not happening.
Figure 1: Average training Categorical-Crossentropy loss logged every $200$ steps (left) and validation perplexity measured every 12 00012\ 000 steps on a language-stratified held out set of documents (right).
We evaluated the resulting model, Racka, in a two-shot setting on selected tasks from the llm-eval-harness Hungarian benchmark and compared its performance against the original Qwen-3 4B base model. The results are presented in Table 3. The preliminary outcomes, as anticipated in our abstract, are modest, but we see great potential in our models future training and downstream applicability. We could also see a drop from the original Qwen3-4B baseline,in the case of ARC (hu), where the original model attained a score of $32.02$.
Table 3: LM-Eval-Harness results for evaluated models (best results in bold)
Dataset (Metric) Qwen3-4B Racka-4B Qwen3-4B-base
Arc_hu (Acc) $0.3202$ $0.3005$ $0.3792$
Arc_hu (Acc_norm) $0.3844$ $0.3450$ $0.4169$
Hellaswag_hu (Acc) $0.3369$ $0.3407$ $0.3610$
Hellaswag_hu (Acc_norm) $0.4095$ $0.4078$ $0.4557$
MMLU_hu (Acc) $0.5427$ $0.5089$ $0.5965$
TruthfulQA_hu_mc1 (Acc) $0.3177$ $0.3047$ $0.3281$
TruthfulQA_hu_mc2 (Acc) $0.5060$ $0.5040$ $0.5045$
GSM8K_hu (Strict-match) $0.6330$ $0.5299$ $0.6398$
GSM8K_hu (Flexible extract) $0.6285$ $0.5329$ $0.6421$
The HuLU benchmark paints a more positive picture (see Table 4). Racka seems to be behind the base model(s) on HuRTE and HuCommitmentBank, but outperforms them on the other four tasks. On HuWNLI and HuCommitmentBank, it even surpasses the current state-of-the-art 8B model and achieves the best result out of all tested models on the former [^5] All numbers are measured on the validation set..
Table 4: Result of the mentioned 3 models on the datasets of the HULU benchmark, values averaged over multiple runs. Best result per benchmark is highlighted with bold.
@@ 5.2 Continual Pretraining Results
Overall Performance $0.711$ $0.382$ $0.685$ $0.751$ $0.502$ $0.7295$ $0.702$ $0.378$ $0.673$ $0.729$ $0.485$ $0.721$
@@
Finally, Table 5 presents the results of evaluating Racka on generation tasks. Beyond the base model and our own, we also compare against the models reported in the original OpenHuEval paper that are close in size. The results show that Racka delivers competitive performance relative to models twice its size and to the corresponding base models.
Table 5: Results on the datasets of the OpenHuEval Benchmark. Best result per benchmark is highlighted with bold.
†: After careful evaluations, we have encountered a problem with the HuWildBench prompting methods and tokenizer compatibility. We will include the revised results in the final paper as well.
Model HuWildBench HuSimpleQA HuProverbRea HuMatchingFIB HuStandardFIB Overall
WBScore Acc Acc (OE) Acc (2CQ) B Acc, Q Acc B Acc Q Acc
Llama-3.1-Instruct-8B $53.62$ $15.2$ $63.35$ $73.48$ $5.74$ $0.72$ $16.64$ $1.08$ $28.72$
Qwen2.5-Instruct-7B $42.01$ $5.22$ $50.48$ $67.05$ $31.88$ $1.08$ $7.43$ 0 $25.64$
Qwen3-4B t.b.d.† $7.3$ $62.47$ $69.25$ $8.1$ $4.14$ $4.68$ $2.15$ t.b.d.†
Qwen3-4B-Base t.b.d.† $5.9$ $41.15$ 0 $42.3$ $5.58$ 0 0 t.b.d.†
Racka-4B t.b.d.† $7.6$ $58.41$ $77.53$ $22.87$ $2.52$ $18.02$ $1.08$ t.b.d.†
@@ 6 Conclusion and Future Directions
This work introduced Racka, a resource-efficient, continually pretrained language model for Hungarian, using a practical methodology specifically designed for deployment on national HPC infrastructure. Through targeted tokenizer adaptation by expanding and initializing the vocabulary to better capture the morphological complexity of Hungarian, we achieved a substantial reduction in subword fertility, improving both computational efficiency and the representational capacity of the model. The construction of a 160B-token multilingual corpus, with a carefully balanced data mixture, further enabled robust language adaptation while reducing catastrophic forgetting of high-resource languages.
The evaluation results partially confirm the effectiveness of our tokenizer adaptation and training strategy. Racka outperforms its base model on some tasks and lags behind it on others. The same relationship can be seen between the reasoning Qwen3-4B and the foundational Qwen3-4B-base models, resulting a complicated picture. Nevertheless, Racka seems to be competitive on most of the benchmarks, sometimes with models twice its size.
Future work includes systematic hyperparameter optimization and supervised fine-tuning for downstream applications as well as preference tuning. We intend to openly publish our model and tokenizer, to support community-driven research and contribute to the advancement of digital sovereignty of Hungary.

### v2

@@ 4.3 Packed Continual Pretraining with LoRA
To enable parameter-efficient continual pretraining, we employed Low-Rank Adaptation (LoRA). LoRA modules were inserted into all linear layers of the Qwen3-4B backbone. The weights of the base model were frozen throughout the training. The embedding and language modeling head layers are tied as in Qwen3-4B.
Our choice of LoRA hyperparameters was guided by recent best practices and empirical validation. Specifically, we set the LoRA rank ($r$) to $64$ and the scaling factor ($\alpha$) to $128$, following the heuristic of setting $\alpha$ to approximately twice the rank to ensure sufficient adaptation capacity. A dropout of $0.1$ was applied to the LoRA adapters to improve generalization. With this setup we train 0.520.52B parameters accounting for $12.5\%$ of the base model size. We used separate learning rates for LoRA and non-LoRA parameters: $1\times 10^{-4}$ for LoRA and $5\times 10^{-5}$ for non-LoRA parameters, both with a weight decay of $0.005$. Optimization was performed using AdamW. The learning rate schedule included a linear warmup phase covering $1\%$ of total training steps, followed by linear decay to zero which is shown to be beneficial (Bergsma et al., 2025). We did not employ a specialized tokenizer healing phase.
To maximize training efficiency, we used input sequence packing (Ding et al., 2024), merging shorter sequences into a single context window of up to 4096 tokens. Both the beginning-of-sequence and end-of-sequence delimiters were set to <|endoftext|>, as defined by the Qwen tokenizer. Packing was performed using a linear scaling algorithm prior to training. For computational efficiency, we used a standard full causal attention mask (as in (DeepSeek-AI, 2025)), rather than a multi-segment lower-triangular mask (Hugging Face, 2024), which would otherwise require quadratic memory to store the mask itself, as well as the inability to use high-speed attention kernels.
@@ 4.4 Training Infrastructure and Parallelization Strategy
We conducted performance measurements on a distributed setup consisting of two nodes, each equipped with four GPUs. In DDP configuration, the training throughput reached 2.92.93.03.0 seconds per iteration. In contrast, FSDP configuration delivered substantially lower performance, reaching 5.05.05.15.1 seconds per iteration. Mixed-precision training with the standard FSDP implementation is not possible, training with full precision resulted in OOM for our minimal training context length of 44k tokens. As DDP yielded superior scaling and GPU utilization, we adopted it for all experiments as a hardware-aware choice for parallelization on the available infrastructure.
@@
Table 2: Tokenizer performance on the validation set: Subword fertility (lower is better) and relative change from Qwen3-4B.
Language Qwen3-4B Racka-4B Change (%)
@@
Table 3: Results on the HULU, OpenHuEval and LM-Eval-Harness benchmarks. HULU results are measured on the publicly available validation set as an average of multiple runs taking the better option from the LoRA and full finetuned models. In case of Qwen and Racka models we patch the original OpenHuEval implementation to be compatible with these models. For details see Appendix C. LM-Eval-Harness was evaluated with and without chat templates and the best overall result is kept (with chat template for Racka and without it for the other models). The best result per benchmark is highlighted with bold.
HULU benchmark
@@
Overall Performance (HULU) $0.711$ $0.382$ $0.685$ $0.751$ $0.502$ $0.7295$ $0.702$ $0.378$ $0.673$ $0.729$ $0.485$ $0.721$
OpenHuEval benchmark
HuWildBench (WBScore) $63.03$ $57.17$ $52.59$ $17.77$
HuSimpleQA (Acc) $7.30$ $10.05$ $5.91$ $20.03$
HuProverbRea (Acc OE) $62.47$ $61.94$ $41.15$ $75.86$
HuProverbRea (Acc 2CQ) $74.98$ $77.53$ 0 $77.36$
HuMatchingFIB (B Acc) $39.59$ $38.93$ $42.3$ $33.54$
HuMatchingFIB (Q Acc) $5.94$ $4.68$ $5.58$ $3.96$
HuStandardFIB (B Acc) $13.20$ $18.98$ 0 $29.16$
HuStandardFIB (Q Acc) $1.08$ $2.15$ 0 $2.15$
Overall Performance (OpenHuEval) $33.44$ 33.928 7533.928\,75 $18.44$ $32.47$
LM-Eval-Harness benchmark
Arc_hu (Acc) $0.3202$ $0.3450$ $0.3792$ $0.3861$
Arc_hu (Acc_norm) $0.3844$ $0.4101$ $0.4169$ $0.4323$
Hellaswag_hu (Acc) $0.3369$ $0.3656$ $0.3610$ $0.4241$
Hellaswag_hu (Acc_norm) $0.4095$ $0.4510$ $0.4557$ $0.5606$
MMLU_hu (Acc) $0.5427$ $0.5378$ $0.5965$ $0.5310$
TruthfulQA_hu_mc1 (Acc) $0.3177$ $0.3644$ $0.3281$ $0.3035$
TruthfulQA_hu_mc2 (Acc) $0.5102$ $0.5493$ $0.5045$ $0.4883$
GSM8K_hu (Strict-match) $0.6330$ $0.5299$ $0.6398$ 0.476 110.476\,11
GSM8K_hu (Flexible extract) $0.6285$ $0.5329$ $0.6421$ $0.4791$
Overall Performance (LM-Eval-Harness) $0.453$ $0.454$ $0.4805$ $0.4546$
4.5 Evaluation Protocol
We assess the model from two perspectives: intrinsic language modeling capabilities and downstream Hungarian benchmark performance.
Intrinsic Evaluation: We use perplexity (PPL) as our primary intrinsic metric. Perplexity measures the model’s uncertainty in predicting the next token in a sequence; a lower perplexity score indicates a better fit to the data distribution (Jelinek et al., 2005). We calculated perplexity on the held-out validation set containing a mix of Hungarian, English, German and code to monitor learning progress and assess for catastrophic forgetting.
LM-Eval-Harness benchmark: To measure the model’s practical utility without task-specific training, we conducted few-shot evaluations on a selection of Hungarian language understanding tasks (Gao et al., 2024). For this we utilized the LM-Eval-Harness with Hungarian translated benchmarks. We used the already implemented benchmarks ARC_HU, MMLU_HU, Hellaswag_HU and TruthfulQA_HU from LM-Eval-Harness, which is the work of (Lai et al., 2023). Additionally, we manually added the Hungarian translated version of GSM8K (Cobbe et al., 2021), which is a mathematical reasoning dataset and was transletd by (Thellmann et al., 2024), to check the model reasoning capabilities on mathemical problems. We ran all benchmarks separately, using the default config parameters with and without chat templates as defined in the original LM-Eval-Harness repository [^4] https://github.com/EleutherAI/lm-evaluation-harness. Table 3 shows our results.
HULU Benchmark The HULU benchmark offers a comprehensive set of supervised tasks covering a wide range of linguistic tasks, including acceptability, causal reasoning, sentiment classification, pronoun resolution, factual commitment, and reading comprehension. For the evaluation, we separately fine-tuned our model on each HULU subtask using a custom fork of the official benchmarking script [^5] https://github.com/nytud/HuLU. Both full and LoRA mode finetuning was performed with the following key parameters: lora_rank=64lora\_rank=64, lora_alpha=128lora\_alpha=128, lora_dropout=0.1lora\_dropout=0.1, and batch_size=2batch\_size=2. We report the best mode as detailed in Appendix D.
OpenHuEval benchmark To further evaluate the ability of the model to perform Hungarian reading comprehension and generation tasks, we conducted fine-tuned downstream evaluations using the OpenHuEval benchmark suite (Yang et al., 2025b). This collection focuses on cloze-style question answering and comprehension tasks designed to capture various aspects of contextual understanding, reasoning, and lexical inference in Hungarian. OpenHuEval comprises five datasets: HuWildBench (real Hungarian user queries), HuSimpleQA (fact-based Hungarian questions), HuProverbRea (reasoning with Hungarian proverbs), HuMatchingFIB (option-based fill-in-the-blank), and HuStandardFIB (free-form fill-in-the-blank). The evaluation was made using the original repository of the benchmark [^6] https://github.com/opendatalab/OpenHuEval.
5 Results and Analysis
To measure the performance of our methodology, in this section we analyze both the tokenizer and the Hungarian performance of Racka-4B. Benchmarks include tasks related to base language use, natural language inference, logic and Hungarian-specific knowledge. We compare our model to the original Qwen3-4B model, the vanilla Qwen3-4B-Base language model and a state-of-the-art Hungarian model, Puli-LlumiX-Llama-3.1 8B. The OTP models are not openly available, benchmarking them is not currently possible. Extending the benchmarks with international models, which according to our experiences tend to perform lower than models trained in Hungary, is matter of future works.
5.1 Tokenizer Performance
The primary goal of our tokenizer adaptation was to improve the encoding efficiency for Hungarian without degrading performance on the high-resource languages present in the base model. We evaluated this using subword fertility, which is the average number of tokens required to represent a single whitespace delimited word (Ács et al., 2021).
The results measured on the validation documents, presented in Table 2, confirm the success of our adaptation strategy. For Hungarian, the Racka tokenizer demonstrates a substantial improvement over the original Qwen3 tokenizer. Subword fertility was reduced by over $46\%$, from $3.13$ down to $1.66$. This means that, on average, the same Hungarian text can be represented with almost half of the original Qwen tokens, leading to direct, in most cases linear improvements in inference efficiency and latency. Crucially, this improvement in Hungarian was achieved with only modest trade-offs for other languages: fertility increased by less than 13% for German and 24% for English. This indicates that our vocabulary extension successfully preserved the efficiency for high-resource languages.
5.2 Continual Pretraining Results
Following the successful tokenizer adaptation, we successfully conducted the continual pretraining phase on the full 160B token dataset. The training ran for 326 357326\,357 steps on $64$ GPUs in parallel, with a per-device batch size of $2$ and gradient accumulation set to $4$, resulting in an effective logical batch size of $512$ sequences with a context length of $4096$ for each model update. The training loss curve is depicted in Appendix E. We see that even though the training loss begins to flatten after 50 00050\,000 steps, the validation perplexity continues to steadily decrease to a final value of $6.99$, indicating that overfitting is not reached.
The HuLU benchmark paints a positive picture (see the top section of Table 3) with a top overall performance. Racka seems to be behind the base model(s) on HuRTE and HuCommitmentBank, but outperforms them on the other four tasks. On HuWNLI and HuCommitmentBank, it even surpasses the current state-of-the-art 8B Puli model and achieves the best result out of all tested models on the former [^7] All numbers are measured on the validation set..
@@ 5.2 Continual Pretraining Results
The middle section of Table 3 presents the results of evaluating Racka on OpenHuEval. We also denote that larger (77-88 B) models in the original OpenHuEval Yang et al. (2025b) paper have an overall score of $28.77$ (LLama-3.1 8B) and $25.64$ (Qwen2.5-instruct 7B). The results show that Racka delivers competitive performance relative to models twice its size and to the corresponding base models with a higher overall score. Racka performs well in tasks that rely on interpretation and reasoning and a lower score on infilling and QA tasks.
Finally, we evaluate Racka on tasks from the LM-Eval-Harness Hungarian benchmark. The results are presented at the bottom of Table 3. On this benchmark the overall performance of Racka is on par with the original Qwen3-4B and the larger Puli model, with Qwen3-4B-Base leading the benchmarks. The results indicate that Racka exhibits reduced mathematical reasoning capabilities compared to the base model, but demonstrates improved performance on question-answering tasks. Notably, Racka outperforms PULI on TruthfulQA, contrasting with its lower performance on HuSimpleQA. We attribute this discrepancy to the broader, less culture-specific nature of TruthfulQA questions, and the limitations of Racka’s culture-specific fact memorization capacity compared to a model twice its size.
@@ 6 Conclusion and Future Directions
This work introduced Racka-4B, a resource-efficient, continually pretrained language model for Hungarian, using a practical methodology specifically designed for deployment on the national Komondor HPC infrastructure. Through targeted tokenizer adaptation by expanding and initializing the vocabulary to better capture Hungarian specific vocabulary elements, we achieved a substantial reduction in subword fertility, improving both computational efficiency and the representational capacity of the model. The construction of a 160B-token multilingual corpus, with a carefully balanced data mixture, further enabled robust language adaptation while reducing catastrophic forgetting of high-resource languages.
The evaluation results confirm the effectiveness of our tokenizer adaptation and training strategy. With the generation speed improvements outlined above, the overall performance of Racka-4B is slightly higher than the base Qwen3-4B models, in some cases topping 8B models such as the latest Puli.
Future work includes systematic hyperparameter optimization, extended benchmarking and supervised fine-tuning for downstream applications as well as preference tuning. We openly publish our model and tokenizer, to support community-driven research and the advancement of digital sovereignty of Hungary.
@@ B Composition of the Hungarian Corpus
The Repositories dataset contains books and electronic journals (academic and otherwise) from the following sources:
EPA
The Elektronikus Periodika Adatbázis [^8] https://epa.oszk.hu/ (Database of Electronic Periodicals)
MEK
OAI
Open Archives Initiative repositories
OJS
Open Journal System repositories
@@
C Fixing OpenHuEval for Qwen3-like Models
We have encountered several difficulties running OpenHuEval. The prompts and configurations contain minor mistakes and reasoning models, such as Qwen3, are not always evaluated correctly. Some of these problems are rooted in the OpenCompass framework itself. The most glaring issue was the omission of the OpenHuEval example scripts assumed to be present by the documentation. This was remedied by the authors after we brought it to their attention.
We have also found that the OpenHuEval tasks handled the reasoning output of models inconsistently. Some, such as HuMatchingFIB, detected the presence of the reasoning tokens and removed them before evaluation; some, such as HuSimpleQA, did not. In many cases, the presence of reasoning traces resulted in a parsing error or the model running out of the allotted context. By default, we wanted to disable reasoning in these benchmarks as small reasoning models are prone to getting stuck in thinking loops when running them with temperature 0, which is needed for reproducibility. This behavior is also outlined in the Qwen3 documentation.
Due to time constraints, we decided to fix these issues locally, with as little change as possible in the OpenCompass codebase: we modified the chat template to respect the enable_thinking tokenizer argument in the model configuration and disabled reasoning for the affected tasks. We made sure to leave the benchmark’s datasets unchanged. The only exception was the HuWildBench task, which posed additional challenges that had to be addressed.
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
    Increasing model generation control. As mentioned above, we make sure that the tokenizer settings are respected to avoid generating reasoning traces. We also enforce repetition penalty and frequency penalty using the definitions of vLLM Kwon et al. (2023), which are both set to $0.3$ in the case of all Qwen and Racka models to avoid repetitive text generation.
  4. 4.
    Fixing judge model errors. HuWildBench is one of the tasks where the original implementation fails to remove (often English) reasoning traces from the judge model prompt. We reimplement this behavior correctly to align with the original intent present in the benchmark configuration. We also find that the judge model rarely follows the JSON schema defined in its prompts. In some cases this results in the OpenCompass implementation missing the judge score. We use a simple regex pattern to find the last “score:” in the ill-formed judge outputs and extract the first integer that follows it as a score. This method is able to correctly handle $99.91\%$ of the failed judge prompts. When no “score:” segment is present in the judge output use a default score of $1$. We enforce the usage of the single model scoring prompt to make sure OpenCompass does not automatically switch to the pairwise evaluation where a parent/base model is present.
While HuWildBench was the only subset where these errors fully prevented the original implementation from working, we believe that future work should revisit the prompts and implementation of OpenHuEval to ensure language quality and compatibility with chat and reasoning language models.
D HULU Downstream Training Details
We average results over multiple trainings to control for random factors during downstream training. The model is trained with early stopping for $10$ fine tuning rounds on HuCOPA, HuWNLI and HuCommitmentBank and $5$ trainings on the other subsets.
We also perform both LoRA and full fine-tuning and report the superior method in Table 3. Here we detail the best method for each model and benchmark subset in Table 4.
Table 4: Details on which fine-tuning method is superior for each model-task combination.
HULU subset Qwen3-4B Racka-4B Qwen3-4B-Base Puli-LlumiX-Llama-3.1 8B
HuCOLA LoRA LoRA LoRA LoRA
HuCOPA LoRA Full LoRA LoRA
HuSST LoRA LoRA LoRA LoRA
HuRTE LoRA LoRA LoRA LoRA
HuWNLI Full Full Full LoRA
HuCommitmentBank Full Full Full LoRA
E Loss Curves
Refer to captionRefer to caption
Figure 1: Average training Categorical-Crossentropy loss logged every $200$ steps (top) and validation perplexity measured every 12 00012\ 000 steps on a language-stratified held out set of documents (bottom).